OpenStudy (anonymous):

Alright guys I'm really stuck on this one.. I'm trying to solve a Log problem but I cant figure out what step is next... 3Ln(X + 4) -5 = 3 First step: add five to both sides 3Ln (X + 4) -5 = 3 +5 +5 3Ln (x + 4) = 8 What do I do next???

OpenStudy (anonymous):

Divide by 3 on both sides

OpenStudy (anonymous):

thn carry on

OpenStudy (anonymous):

so it should look something like this then, right... 3 Ln ( X + 4) = 8 ---- --- 3 3 Can you walk me through each step, please? I would really appreciate it

OpenStudy (anonymous):

Yes.. I will assist you... now what will be the equation?

OpenStudy (anonymous):

My question is to walk me through the next couple of steps after I divide by 3 on both sides. Can you show me what it should look like? Thanks..

OpenStudy (anonymous):

mmm.. you mean I have to write each steps??

OpenStudy (anonymous):

well if you could just briefly explain to me what to do after I divide by 3 that would be helpful.... I'm just completely lost on this problem. I pretty much only know how to do the first step

OpenStudy (anonymous):

Ok.. afet dividing by 3 you have to take inverse logarithm

OpenStudy (anonymous):

then you will be abl to find the value of x

OpenStudy (anonymous):

so what might that look like... could you write out the steps by placing the numbers in the correct spots. I know I'm asking a lot but it would really help me understand the problem better... :/

OpenStudy (anonymous):

actually writting step by step or direct answr is against the CoC... So I will write an example

zepdrix (zepdrix):

Exponentiation is the inverse of the logarithmic function. Here's a quick example. \[\large \ln x = 2\] This is a log with base e, we'll exponentiate both sides, rewrite both sides as exponents with base e. \[\huge e^{\ln x}=e^2\] Since the exponential and the logarithm are inverse operations of one another, they essentially "cancel out". \[\huge x=e^2\]

OpenStudy (anonymous):

well examole by @zepdrix is sufficient , I guess... You have to work it out

OpenStudy (anonymous):

Ohhh ok, so natural Log (Ln) and (e) cancel out. ok. Well that does help out a little bit. Thank you for your time.

OpenStudy (anonymous):

welcome......

zepdrix (zepdrix):

Yah performing the inverse operation of the logarithm is a bit tricky to get used to :) If you're still confused on how to perform that on your particular problem, just let us know.

OpenStudy (anonymous):

Thank you, I'm sure i'll figure it out. Your explanation did help. Yes, Inverse operation of Logarithm is tricky.

OpenStudy (anonymous):

all the best....